OpenStudy (anonymous):

Set A consists of the ten digits 0,0,0,0,0,0,2,2,2,4. Set B consists of the seven digits 0,0,0,0,2,2,2. One digit is chosen at random from each set. The random variable X is defined as the sum of these two digits. (i) Show that P(X=2)=3/7 (ii)Tabulate the probability distribution X (iii)Find E(X) and Var(X) (iv)Given that X=2, find the probability that the digit chosen from set A was 2

OpenStudy (anonymous):

please help... I'm confused here 0_0

OpenStudy (anonymous):

Need help?

OpenStudy (anonymous):

First, let's start with "i". And, for simplicity, lets say that choosing from set A is signified by the random variable "A" (and "B" for set B). That means that X = A+B Well, "i" is simple because there are only two ways that X can equal "2": Either A=0 and B=2 Or A=2 and B=0 So calculate the probability of these two events and add them together: \[\frac{ 3 }{ 5 }*\frac{ 3 }{ 7 }+\frac{ 3 }{ 10 }*\frac{ 4 }{ 7 } = \frac{ 3 }{ 7 }\] With me so far?

OpenStudy (anonymous):

yes :D now i understand where you get the values